Traffic signs are regulated by the Manual on Uniform Traffic Control Devices. The perimeter of some traffic signs must be 80 inches. Also its length is 8 inches longer than its width. Find the dimensions of this sign.
Answer by asuar010(338)

the perimeter formula is 2(l+w)=80 so l+w=40 and if we substitute we get that the length is 8+w so we get 8+w+w=40 so 2w=32 so w=16 and l=24.